Solve the following system:
x - 2y = 5
3x + 2y = -1
What is the value of x?

Found 2 solutions by jim_thompson5910, amarjeeth123:
Answer by jim_thompson5910(35256)

Add up the two equations to get

x - 2y = 5
3x + 2y = -1
-------------
4x + 0y = 4

So if 4x = 4, then x = 1 (divide both sides by 4 to isolate x)

Answer by amarjeeth123(569)

x - 2y = 5...........equation 1
3x + 2y = -1.........equation 2
Adding 1 and 2 we get,
x+3x=5-1
4x=4
x=1
y=-2